A built-up beam has the following cross section: 7.5 cm 7.5 cm 10 cm 5 cm 4 cm 6 cm where x = 5 cm
Answer #1

Moment of inertia of piece 2 about x axis;

Ix_2 = bh3/12 + A2d22

b = 15 cm

h = 5 cm

A2 = b*h = 15*5 = 75

d2 = 5/2 = 2.5 cm

Ix_2 = 15*5*5*5/12 + 75*2.5*2.5 = 156.25 + 468.75 = 625 cm4